the sum of the digits of a two-digit number is 14. if the number represented by reversing the digits is subtracted from the origional number, the result is 18. what is the origional number?
Answer by stanbon(48559)

the sum of the digits of a two-digit number is 14.
---
if the number represented by reversing the digits is subtracted from the original number, the result is 18.
---
what is the origional number?
==================================
Let the original number be 10t+u
Equations:
t + u = 14
10t+u - (10u+t) = 18
-----------------------------
Rearrange:
t + u = 14
9t-9u = 18
----------------------
Simplify:
t+u = 14
t-u = 2
--------------
Add and solve for "t"
2t = 16
t = 8
--------------
Substitute into t+u = 14 to get u = 6
==========================================
Original Number: 86
==========================================
Cheers,
Stan H.